OpenStudy (griffindore):

What is the value of the expression 3.6 x 10^8 over 1.2 x 10^3? 2.4 ⋅ 10^5 3.0 ⋅ 10^11 3.0 ⋅ 10^5 2.4 ⋅ 10^11

OpenStudy (welshfella):

3.6 / 1.2 = ?

OpenStudy (welshfella):

10^8 / 10^3 = 10^(8-3)

OpenStudy (welshfella):

can you complete those 2 calculations?